# Low-Altitude Wireless Networks: A Survey

### v1

@@ 3.3 Networked Control
Networked control serves as a cornerstone of distributed cyber-physical systems, enabling the seamless integration of communication and control functions through shared network infrastructures [73, 74]. By facilitating real-time exchange of sensory information and control commands among distributed nodes, it provides a scalable mechanism for coordinated operation and decision-making.
@@
This evolutionary trajectory highlights how LAWNs transform rigid and isolated UAV control architectures into intelligent, resilient, and cooperative swarm systems. Such advancements are instrumental in supporting the LAE, where networked UAV swarms serve as the backbone for mission-critical LAE applications including urban surveillance, dynamic aerial coverage, and logistics distribution [75], all of which require precise synchronization and adaptive coordination in complex low-altitude environments. Through real-time sharing of control signals and sensory information, each UAV dynamically adjusts its trajectory and task allocation based on both local observations and global swarm feedback, thereby ensuring efficient, safe, and scalable operations across diverse LAE scenarios.
@@
This conceptual evolution is further supported by empirical research, where LAWNs-enabled networked control has been shown to improve robustness and adaptability in real-world conditions, providing essential technical support for reliable operations in LAE scenarios. For instance, [76] proposed a multi-UAV cooperative framework specifically designed for dynamic electromagnetic environments, leveraging UAV mobility, adaptive coverage, and reliable line-of-sight connectivity to enhance perception and overall system robustness. Complementing this approach, closed-loop feedback mechanisms were shown to significantly improve swarm adaptability under interference, uncertainty, and partial node failures [77]. Furthermore, [78] introduced a distributed optimal control framework for multi-UAV systems that simultaneously supported formation management, obstacle avoidance, and connectivity maintenance while adaptively optimizing communication resources.
@@
Beyond robustness and adaptability, recent studies highlighted the benefits of joint optimization across sensing, communication, and control to achieve end-to-end quality of service (QoS), a capability that can directly enhance mission reliability in critical LAE applications. For instance, [79] proposed a joint optimization framework for UAV-based positioning, where UAVs acted as aerial anchors to ensure reliable localization in challenging environments. This approach reduced service failure rates by over $75\%$ and resource consumption by $80\%$, demonstrating the efficiency of cross-domain integration. Air-ground collaboration was also explored in [80], where a formation control strategy incorporating spatial synergy enabled UAVs to provide robust relay communication while maintaining coordinated positioning. Likewise, [81] presented a lightweight, UAV-based emergency communication framework employing commercial micro-UAVs. By integrating body-conformal omnidirectional antennas with optimized formation strategies, the system enhanced cooperative transmission and ensured reliable coverage in harsh terrains such as mountainous canyons.
@@
In summary, networked control in LAWNs has evolved from isolated point-to-point architectures to intelligent, distributed, and adaptive control infrastructures. By integrating ISAC-enabled perception, edge intelligence, and resilient communication links, LAWNs provide unprecedented scalability, robustness, and autonomy for UAV swarms. This progression lays the foundation for future networked swarm intelligence. It is capable of operating reliably in dynamic, uncertain, and resource-constrained low-altitude environments, which are increasingly relevant for crowded urban air mobility, large-scale event management, and disaster-stricken regions.
@@ 3.4 Wireless Power Transfer
Wireless power transfer (WPT) delivers energy to electronic devices through radio-frequency electromagnetic waves or inductive coupling, thereby enabling continuous power replenishment without physical connections [82]. In LAWNs, WPT provides an effective solution to the endurance limitations of battery-powered UAVs, extending mission duration, reducing task interruptions, and minimizing maintenance costs [83]. Its integration with UAV swarms further enhances operational sustainability, making long-duration and high-frequency missions feasible. The evolution of WPT in UAV-enabled networks can be conceptually divided into three progressive stages, as illustrated in Fig. 4.
@@
Extending the concept of cooperative energy sharing, recent research has explored the implementation of LAWNs-enabled large-scale WPT in practical scenarios, which can serve as a key technological enabler for various LAE applications. In [84], a cooperative wireless-powered MEC framework was proposed, where a grid-powered access point provided both computational services and laser-based charging to a UAV. The UAV, in turn, functioned as an information relay, energy relay, and mobile edge server. Similarly, [85] introduced a UAV-assisted cooperative system that employed simultaneous wireless information and power transfer (SWIPT), enabling UAVs to act as mobile relays powered directly by RF energy. Beyond cooperative relaying, heterogeneous architectures were explored. For instance, [86] proposed a UAV-enabled MEC framework where UAVs served as both wireless chargers and mobile computing enhancers for energy-constrained IoT devices. Likewise, [87] considered a harvest-then-transmit protocol in UAV-enabled wireless powered communication networks, while [88] investigated trajectory optimization in obstacle-rich environments to maximize energy delivery efficiency. At the ground-user level, [89] optimized 3D UAV trajectories for fair and efficient power allocation among distributed nodes, and [90] addressed charging of energy-constrained sensor nodes with unknown locations via UAV-based WPT strategies.
In summary, WPT in LAWNs evolves from static ground-based charging to a dynamic, network-wide energy sharing paradigm. By leveraging UAV mobility, swarm cooperation, and joint optimization with sensing and communication, LAWNs enable persistent, autonomous, and large-scale aerial operations. When combined with ISAC and edge intelligence, WPT does not merely extend flight time. It provides the energy backbone that sustains intelligent perception, resilient communication, and adaptive networked control for low-altitude economic scenarios such as long-endurance surveillance, infrastructure inspection, cargo transport, and emergency energy replenishment.
4 LAWN Performance Metrics
In this section, we survey how key performance metrics in terms of latency, energy efficiency (EE), and scalability& collectivity have been addressed in LAWNs, focusing on how they guide concrete design trade-offs and optimization frameworks.
4.1 Timeliness and Latency
In practice, LAWNs are frequently deployed to support time-sensitive applications, where the transmission delay may lead to severe operational and safety implications. To name a few, in emergency response, the timely delivery of situational data such as survivor locations, fire spread, or hazardous material detection is essential for enabling rapid decision-making and coordinated rescue actions. Therefore, one of the core design principles for LAWNs is the minimization of end-to-end latency while maintaining reliable data transmission. To achieve this, finite-block length (FBL) transmission has emerged as an important technique for managing communication in constrained networks. Unlike the traditional Shannon capacity framework, which assumes infinitely long codewords, FBL transmission provides a more practical characterization of the achievable rates given block lengths and error probability. More importantly, the channel conditions in LAWNs often change rapidly, leading to the reliable infinite block length transmission assumption being invalid. The authors of [91, 92, 93] investigated finite blocklength multiterminal lossy source coding, providing new benchmarks into efficient compression with multiple encoders, decoders, and source sequences, thereby strengthening the theoretical foundation of multiterminal low-latency communications and greatly promoting the development of low-altitude low-latency networking. The work of [94] investigated IoT service provisioning by leveraging FBL transmission in reducing communication delays within drone-aided MEC networks, with drones being both communication relays and computational resources. Jin et al. proposed the co-design of sensing, communication, and control in LAWNs, where FBL transmission is utilized with remote control tasks to minimize latency. The results show that optimizing resource allocation for communication and control can lead to lower delays and better overall performance [95].
Some recent studies focus on low-latency communications. For instance, the authors in [96] presented a predictive precoder design for ultra-reliable low-latency communication (URLLC) systems, where a deep learning (DL)-based framework was developed to predict the precoder via historical channel data. By leveraging a convolutional long short-term memory (CLSTM) network, reliable communication with minimal latency can be guaranteed, which demonstrates superior frame error rate (FER) performance approaching the theoretical lower bound with reduced latency. More relatively, the work of [97] explored resource allocation for a two-way UAV relaying system under URLLC constraints, focusing on minimizing latency in both the forward and backward links. An iterative approach was proposed to maximize the system throughput while fulfilling the latency and decoding error probability requirements.
Beyond FBL transmission in the physical layer, the AoI is another critical metric facilitating delay-aware LAWN design. In particular, the AoI measures the freshness of data in a communication network, where outdated data can lead to poor decision-making performance. Toward this end, the study in [98] explored UAV-assisted wireless-powered IoT networks, which optimized UAV trajectories to minimize AoI and ensure that the sensor data remains fresh. Furthermore, Gao et al. examined AoI-sensitive data collection in multi-UAV-assisted wireless sensor networks [99], in which a clustering algorithm, devised to optimize both AoI and energy consumption in multi-UAV systems, was developed. The results show that by adjusting the UAV flight paths and clustering strategies, AoI can be minimized, such that LAWN nodes have access to the freshest data with low energy usage, highlighting that AoI management is essential for real-time collaboration and successful task execution within LAWNs. The importance of real-time scheduling in LAWN systems was studied in [100] to manage communication effectively. By prioritizing the transmission of high-priority data and optimizing the transmission schedules, real-time scheduling can help minimize latency and AoI.
4.2 Energy-Efficient Design
The sustainability of LAWNs are inherently constrained by the limited onboard energy and computational resources of UAVs, making energy-efficient design essential for long-term viability in real-world applications. Early studies primarily examined energy-aware deployment strategies to extend system lifetime. For instance, [101] proposed energy-saving deployment algorithms for UAV swarms to optimize wireless coverage while minimizing energy consumption. In a similar approach, the study of [102] addressed the rapid deployment problem of heterogeneous UAVs to provide wireless coverage over a target area, where the objective for minimizing the maximum deployment delay to ensure fairness and minimizing the total deployment delay to enhance efficiency were addressed, respectively. However, these works largely focus on stationary optimization and do not fully exploit the high-mobility characteristics of LAWNs. In the context, the authors in [103] investigated joint trajectory and transmit power control for UAV relay networks to minimize the outage probability. By carefully devising UAV trajectories and power allocation, energy consumption can be reduced without sacrificing communication reliability. Further, the study in [104] proposed a UAV-assisted backscatter communication system for energy-efficient IoT, which shows a tradeoff between collection location and outage probability on the improvement of EE. On this basis, the work of [105] investigated the trade-off between communication/sensing performance and energy consumption in ISAC-LAWN systems, where the goal was to maximize worst-case EE by jointly designing beamforming and UAV trajectory.
Note that [101, 102, 103, 104, 105] primarily focused on the transmit power control, significantly lower than the propulsion-related power. Hence, the work of [106] proposed a LAWN system simultaneously sensing ground users and forwarding the sensed information. The authors formulated a multi-objective optimization problem to maximize the EE by jointly optimizing user scheduling, movement-related power, and UAV trajectory. While existing methods have improved network EE, their optimization objectives are often limited to reducing the energy consumption of individual UAVs, overlooking the potential gains from cooperative flight dynamics in LAWNs. In real-world deployments, UAVs inevitably interact with surrounding airflow patterns, which can be strategically exploited to boost overall EE. Specifically, airflow over UAV wings generates wake vortices that create an upwash region behind the wingtips. A trailing UAV entering this region experiences reduced aerodynamic drag, thereby lowering propulsion energy consumption, as well-documented in avian flocking. Building on this principle, [107, 108] introduced an ISAC-LAWN design that jointly takes into account aerodynamic factors and sensing/communication performance, showing that a ’V’-shaped formation is likely to be aerodynamically favorable.
It is evident that the above approaches can only save energy to extend network lifetime, lacking the ability to recharge the battery to achieve a long-term power supply. To address this challenge, the authors in [109] explored the performance improvement of UAV communication systems integrated with solar energy harvesting. The uncertainty of solar energy supply was investigated, followed by presenting an iterative optimization approach to maximize throughput.. The simulations show that integrating solar energy into LAWN systems can significantly extend the operational duration, thereby enhancing the sustainability in remote areas where recharging infrastructure is scarce. Meanwhile, the work of [110] investigated energy-efficient LAWN communications design with energy harvesting, where trajectory optimization techniques were used to minimize energy consumption while maximizing the coverage area as well as maintaining QoS.
4.3 Scalability and Connectivity
Benefiting from the high mobility and on-demand deployment capabilities, scalability and connectivity are essential for LAWNs to adapt to various practical requirements. The performance of large-scale LAWN networks is heavily influenced by the ability to manage communication capacity, coverage area, and resilience under dynamic conditions, such as UAV movement and interference. For instance, to reveal the fundamental performance limits of beam alignment for mobile targets such as UAVs in low-altitude networks, the authors in [111, 112] investigated mmWave multiple-antenna communications, deriving theoretical benchmarks for optimal non-adaptive query procedures in multidimensional target search and extending them to scenarios with multiple and moving targets. They further demonstrated the benefits of adaptive query procedures, thereby providing finite-blocklength benchmarks that enable reliable connectivity in low-altitude networks. As evidenced by [113], wherein network localization and formation control in LAWN systems with asynchronous agents were investigated. Specifically, the authors developed a message-passing algorithm using Gaussian messages for efficient localization and clock offset estimation to maintain robust information exchange across all LAWN nodes. Subsequently, two types of control policies, namely angle-based and distance-based formation control, were employed for adaptive obstacle avoidance. By doing so, the LAWN systems are capable of quickly responding to environmental changes, and thereby, enhancing scalability and connectivity.
However, the coexistence of heterogeneous LAWN nodes sharing spectral and spatial domains creates a complex interference environment that challenges large-scale deployment and reliable interconnection. The main challenges stem from beam pattern overlaps and waveform interactions across both spatial and temporal dimensions. To tackle these issues, the work in [117] introduced a mode-switching strategy, where UAVs alternate between jamming and information transmission. By optimizing mode selection, scheduling, and trajectories, this approach ensures strong A2G links even in the presence of mobile ground users. Furthermore, [118] proposed a deep learning–based predictive beamforming scheme that mitigated UAV jittering by forecasting the angles between UAV and user, thereby enabling robust and reliable UAV communications. Moreover, the authors in [119] proposed a dynamic interference management framework by jointly optimizing UAV trajectories and power allocation, thereby improving overall network efficiency under dense deployments. Beyond intra-LAWN coordination, such solutions can also be extended to satellite integration. For example, the authors in [120] designed robust beamforming schemes for space–air–ground integrated networks (SAGIN) to guarantee reliable downlink communications.
@@ 5.1 Privacy preservation
Privacy preservation of LAWNs refers to protecting sensitive information and preventing unauthorized access or disclosure of data related to the operation, environment, and participants within LAWNs [12, 121]. By addressing the privacy concerns of the above issues, LAWNs can foster user trust, enable wider adoption, and comply with legal and ethical requirements, ultimately unlocking the full potential of LAWN technology while minimizing the risks associated with data breaches and misuse. Currently, researchers have adopted differential privacy, federated learning, homomorphic encryption, secure multi-party computation, destination obfuscation, game theory and blockchain-based techniques to realize the privacy preservation of LAWNs on sensing, communication, computing, and control aspects.
Figure 5: The framework of the proposed SEAL defense scheme in UAV offloading scenario [122].
@@
For example, for privacy preservation on the sensing aspect, to ensure privacy in UAV-based delivery and sensing, [123] introduced a game-theoretic routing framework that aims to protect UAV trajectory privacy. The approach leverages a task-time graph and a non-cooperative potential game to optimize sensing, routing, and task selection. This framework demonstrates effective privacy preservation while ensuring timely delivery and achieving efficiency. In terms of communication privacy, [124] presented a traceable and privacy-preserving authentication scheme for UAV applications, employing digital signatures to ensure integrity, confidentiality, and manage sensitive area access. Addressing security in UAV surveillance, [125] introduced a smart Internet of drone (S-IoD) framework with IPA and a lightweight privacy-preserving scheme (L-PPS). This L-PPS expedites authentication between devices through a secret token and dynamic user authentication, mitigating computational overhead. For privacy in UAV-assisted IoT computation offloading, [126] identified that current deep reinforcement learning-based methods leak UAV offloading preferences. The authors introduced a privacy-preserving method that integrates differential privacy into deep reinforcement learning to protect these preferences, including formal analysis of security and utility. Similarly, as shown in Fig. 5, [122] addressed the critical requirements of incentives, privacy, and fairness when offloading UAV computation tasks to ground vehicles. Their proposed framework overcomes issues of manipulation and privacy leakage by using a strategy-proof reverse combinatorial auction. In addition, fairness is managed through smart contracts and hash-chain micropayments, complemented by a privacy-preserving off-chain auction protocol utilizing a trusted processor. Focusing on control-related privacy, [127] addressed the privacy-preserving navigation (PPN) problem for visual-based UAVs, aiming to minimize privacy violations within specific sub-areas while covering a target region through a heuristic path planning solution. The work of [128] tackled privacy concerns in cooperative UAV operations by proposing a decentralized framework that preserves trajectory and destination privacy through probabilistic destination obfuscation and a noncooperative Bayesian game for joint routing and charging.

### v2

@@ 4.2 Networked Control
Networked control serves as a cornerstone of distributed cyber-physical systems, enabling the seamless integration of communication and control functions through shared network infrastructures [meng2024communication, jin2025predictive]. By facilitating real-time exchange of sensory information and control commands among distributed nodes, it provides a scalable mechanism for coordinated operation and decision-making.
@@
This evolutionary trajectory highlights how LAWNs transform rigid and isolated UAV control architectures into intelligent, resilient, and cooperative swarm systems. Such advancements are instrumental in supporting the LAE, where networked UAV swarms serve as the backbone for mission-critical LAE applications including urban surveillance, dynamic aerial coverage, and logistics distribution [mozaffari2019tutorial], all of which require precise synchronization and adaptive coordination in complex low-altitude environments. Through real-time sharing of control signals and sensory information, each UAV dynamically adjusts its trajectory and task allocation based on both local observations and global swarm feedback, thereby ensuring efficient, safe, and scalable operations across diverse LAE scenarios.
@@
This conceptual evolution is further supported by empirical research, where LAWNs-enabled networked control has been shown to improve robustness and adaptability in real-world conditions, providing essential technical support for reliable operations in LAE scenarios. For instance, [chen2025efficient] proposed a multi-UAV cooperative framework specifically designed for dynamic electromagnetic environments, leveraging UAV mobility, adaptive coverage, and reliable line-of-sight connectivity to enhance perception and overall system robustness. Complementing this approach, closed-loop feedback mechanisms were shown to significantly improve swarm adaptability under interference, uncertainty, and partial node failures [qiao2022communication]. Furthermore, [ping2024communication] introduced a distributed optimal control framework for multi-UAV systems that simultaneously supported formation management, obstacle avoidance, and connectivity maintenance while adaptively optimizing communication resources.
@@
Beyond robustness and adaptability, recent studies highlighted the benefits of joint optimization across sensing, communication, and control to achieve end-to-end quality of service (QoS), a capability that can directly enhance mission reliability in critical LAE applications. For instance, [wang2023qos] proposed a joint optimization framework for UAV-based positioning, where UAVs acted as aerial anchors to ensure reliable localization in challenging environments. This approach reduced service failure rates by over $75\%$ and resource consumption by $80\%$, demonstrating the efficiency of cross-domain integration. Air-ground collaboration was also explored in [chen2022formation], where a formation control strategy incorporating spatial synergy enabled UAVs to provide robust relay communication while maintaining coordinated positioning. Likewise, [zhang2020unmanned] presented a lightweight, UAV-based emergency communication framework employing commercial micro-UAVs. By integrating body-conformal omnidirectional antennas with optimized formation strategies, the system enhanced cooperative transmission and ensured reliable coverage in harsh terrains such as mountainous canyons.
@@
In summary, networked control in LAWNs has evolved from isolated point-to-point architectures to intelligent, distributed, and adaptive control infrastructures. This progression lays the foundation for future networked swarm intelligence. It is capable of operating reliably in dynamic, uncertain, and resource-constrained low-altitude environments, which are increasingly relevant for crowded urban air mobility, large-scale event management, and disaster-stricken regions.
4.3 Mobile Edge Computing
Mobile edge computing (MEC) seamlessly integrates computational resources into the network edge [li2025edge], enabling data processing and inference to occur locally at edge nodes such as UAVs, base stations, and edge servers. By minimizing the need to transmit large volumes of raw data to distant cloud data centers, this approach effectively mitigates latency and alleviates bandwidth congestion, thereby enhancing system efficiency, scalability, and resilience.
In the context of LAWNs, the adoption of MEC significantly elevates system performance by exploiting the distributed, mobile, and cooperative nature of UAVs. As depicted in Fig. 5, the evolution of MEC within UAV-enabled networks can be conceptualized in three progressive stages:
  • Stage 1 (Local computing): In this initial stage, each UAV independently executes computational tasks using its limited onboard processing capabilities. While this setup is suitable for lightweight tasks, the constrained computational resources and energy limitations often result in delays when executing more complex algorithms, thereby impeding real-time decision-making and autonomous operations.
  • Stage 2 (Edge-enabled collaborative computing): At this stage, UAVs offload computational tasks to nearby static edge resources, such as base stations or terrestrial edge servers. This enables faster response times and more efficient resource utilization compared to local computing; however, the fixed nature of the edge nodes still limits flexibility in highly dynamic environments.
  • Stage 3 (LAWN-enabled MEC): In this advanced stage, UAVs evolve into dynamic, mobile computing nodes that collectively form a self-organizing aerial-ground edge network. Unlike Stage 2, where edge resources are fixed, UAVs in this stage are capable of mobilizing and coordinating as needed, dynamically deploying additional communication and computational capacity in response to hotspots or mission-critical requirements. Computation is performed adaptively through federated learning, proactive task offloading, and mobility-aware resource orchestration. This stage marks a paradigm shift from static edge collaboration to a flexible, network-wide computing infrastructure empowered by UAV mobility.
This progression is especially pivotal for enabling and sustaining LAE activities, where UAV operations must continuously adapt to dynamic and uncertain environments. LAE-relevant tasks, such as real-time logistics routing, aerial infrastructure inspection, target tracking, and environmental perception, require low-latency and highly reliable computation. Traditional cloud-centric computing models are often constrained by high latency due to long-distance backhaul links, which can hinder time-sensitive applications. In contrast, MEC enables lightweight computational models to be deployed directly on UAVs or adjacent edge nodes, facilitating immediate decision-making and significantly enhancing responsiveness, autonomy, and mission reliability [huynh2022uav]. This results in more efficient and reliable execution across a diverse array of low-altitude economic scenarios.
Furthermore, the distributed nature of LAWNs, with multiple UAVs operating cooperatively, lends itself well to distributed computing paradigms such as federated learning. These paradigms offer scalable, privacy-preserving, and resource-efficient solutions for LAE operations. By training models collaboratively across UAV nodes without sharing raw data, federated learning not only ensures privacy but also reduces communication overhead [zhang2025latency]. For example, [deng2022uav] introduced a UAV-enabled MEC framework for air-ground integrated networks, where UAVs acted as aerial base stations to provide adaptive coverage and edge computing support for Internet-of-Things (IoT) devices under stringent energy and accuracy constraints. Similarly, [wang2023uav] proposed a two-tier hierarchical federated learning framework, where UAVs served as mobile relays, enhancing connectivity to a central server while simultaneously conducting local computations, thus improving scalability in large-scale, infrastructure-limited environments.
Beyond distributed learning, recent advancements in UAV-enabled MEC have explored a variety of application domains and performance metrics directly applicable to LAE services. For instance, [zhang2025age] introduced a UAV-assisted IoT data collection framework under segmented fading channels, which achieved reduced AoI, lower energy consumption, and faster mission completion times through collaborative UAV operations. In [shaikh2025flying], a flying-edge computing framework for smart agriculture was proposed, where UAVs equipped with AI-powered sensors perform tasks such as disease detection and irrigation control on-board, enabling real-time decision-making in areas with limited connectivity. Additionally, [zhang2024minimizing] presented a UAV-assisted MEC framework that optimizes UAV deployment and computation offloading to minimize response delays, ensuring rapid adaptability to dynamic operational environments.
@@ 4.4 Wireless Power Transfer
Wireless power transfer (WPT) delivers energy to electronic devices through radio-frequency electromagnetic waves or inductive coupling, thereby enabling continuous power replenishment without physical connections [zhang2018wireless]. In LAWNs, WPT provides an effective solution to the endurance limitations of battery-powered UAVs, extending mission duration, reducing task interruptions, and minimizing maintenance costs [xu2018uav]. Its integration with UAV swarms further enhances operational sustainability, making long-duration and high-frequency missions feasible. The evolution of WPT in UAV-enabled networks can be conceptually divided into three progressive stages, as illustrated in Fig. 5.
@@
Extending the concept of cooperative energy sharing, recent research has explored the implementation of LAWNs-enabled large-scale WPT in practical scenarios, which can serve as a key technological enabler for various LAE applications. In [hu2020wireless], a cooperative wireless-powered MEC framework was proposed, where a grid-powered access point provided both computational services and laser-based charging to a UAV. The UAV, in turn, functioned as an information relay, energy relay, and mobile edge server. Similarly, [yin2019uav] introduced a UAV-assisted cooperative system that employed simultaneous wireless information and power transfer (SWIPT), enabling UAVs to act as mobile relays powered directly by RF energy. Beyond cooperative relaying, heterogeneous architectures were explored. For instance, [liu2024joint] proposed a UAV-enabled MEC framework where UAVs served as both wireless chargers and mobile computing enhancers for energy-constrained IoT devices. Likewise, [xie2018throughput] considered a harvest-then-transmit protocol in UAV-enabled wireless powered communication networks, while [wang2025practical] investigated trajectory optimization in obstacle-rich environments to maximize energy delivery efficiency. At the ground-user level, [li2023energy] optimized 3D UAV trajectories for fair and efficient power allocation among distributed nodes, and [shi2023two] addressed charging of energy-constrained sensor nodes with unknown locations via UAV-based WPT strategies. In summary, WPT in LAWNs evolves from static ground-based charging to a dynamic, network-wide energy sharing paradigm. By leveraging UAV mobility, swarm cooperation, and joint optimization with sensing and communication, LAWNs enable persistent, autonomous, and large-scale aerial operations. When combined with ISAC and edge intelligence, WPT does not merely extend flight time. It provides the energy backbone that sustains intelligent perception, resilient communication, and adaptive networked control for low-altitude economic scenarios such as long-endurance surveillance, infrastructure inspection, cargo transport, and emergency energy replenishment.
From a system perspective, multi-functional LAWN designs are best understood through task integration rather than as a collection of isolated functions. A concrete mission may simultaneously impose requirements on sensing accuracy, communication reliability, computation offloading, control stability, and energy sustainability. Effective LAWN operation accordingly relies on a coordinated mechanism that maps mission objectives into coupled decisions on beamforming, scheduling, computation placement, controller updates, and power management across heterogeneous aerial and ground nodes. For instance, ISAC determines the timeliness and quality of state and environment information available to edge intelligence and networked control, while control and trajectory decisions in turn reshape channel conditions, interference patterns, and energy consumption. Computation offloading decides how quickly complex functions such as online planning and collaborative perception can be updated and executed in a closed loop, subject to latency and reliability constraints imposed by the communication layer. Meanwhile, WPT further constrains the long-term energy budget of aerial platforms and creates trade-offs between aggressive functional performance and sustainable operation. In practice, multi-functional LAWN design must therefore balance the competing demands of ISAC, edge intelligence, networked control, and WPT within a unified task-centric framework, rather than optimizing each function independently.
@@ 5.1 Privacy preservation
Privacy preservation of LAWNs refers to protecting sensitive information and preventing unauthorized access or disclosure of data related to the operation, environment, and participants within LAWNs [cordill2025comprehensive, kaleem2025quantum]. By addressing the privacy concerns of the above issues, LAWNs can foster user trust, enable wider adoption, and comply with legal and ethical requirements, ultimately unlocking the full potential of LAWN technology while minimizing the risks associated with data breaches and misuse. Currently, researchers have adopted differential privacy, federated learning, homomorphic encryption, secure multi-party computation, destination obfuscation, game theory and blockchain-based techniques to realize the privacy preservation of LAWNs on sensing, communication, computing, and control aspects.
@@
For example, for privacy preservation on the sensing aspect, to ensure privacy in UAV-based delivery and sensing, [liu2023decentralized] introduced a game-theoretic routing framework that aims to protect UAV trajectory privacy. The approach leverages a task-time graph and a non-cooperative potential game to optimize sensing, routing, and task selection. This framework demonstrates effective privacy preservation while ensuring timely delivery and achieving efficiency. In terms of communication privacy, [chen2020traceable] presented a traceable and privacy-preserving authentication scheme for UAV applications, employing digital signatures to ensure integrity, confidentiality, and manage sensitive area access. Addressing security in UAV surveillance, [deebak2020smart] introduced a smart Internet of drone (S-IoD) framework with IPA and a lightweight privacy-preserving scheme (L-PPS). This L-PPS expedites authentication between devices through a secret token and dynamic user authentication, mitigating computational overhead. For privacy in UAV-assisted IoT computation offloading, [wei2021uav] identified that current deep reinforcement learning-based methods leak UAV offloading preferences. The authors introduced a privacy-preserving method that integrates differential privacy into deep reinforcement learning to protect these preferences, including formal analysis of security and utility. Similarly, as shown in Fig. 6, [wang2023seal] addressed the critical requirements of incentives, privacy, and fairness when offloading UAV computation tasks to ground vehicles. Their proposed framework overcomes issues of manipulation and privacy leakage by using a strategy-proof reverse combinatorial auction. In addition, fairness is managed through smart contracts and hash-chain micropayments, complemented by a privacy-preserving off-chain auction protocol utilizing a trusted processor. Focusing on control-related privacy, [pan2019unmanned] addressed the privacy-preserving navigation (PPN) problem for visual-based UAVs, aiming to minimize privacy violations within specific sub-areas while covering a target region through a heuristic path planning solution. The work of [liu2024privacy] tackled privacy concerns in cooperative UAV operations by proposing a decentralized framework that preserves trajectory and destination privacy through probabilistic destination obfuscation and a noncooperative Bayesian game for joint routing and charging.
Figure 6: The framework of the proposed SEAL defense scheme in UAV offloading scenario [wang2023seal].